OpenStudy (icalibear):

Which is the equation for the line perpendicular to y = -5/3x + 11 1/3 and containing P(-2, 3)?

OpenStudy (anonymous):

Have you learned flip and switch?

OpenStudy (icalibear):

No, what is it?

OpenStudy (anonymous):

What is the slope of your equation?

OpenStudy (icalibear):

-5/3

OpenStudy (anonymous):

Correct so a line perpendicular to that function would have a slope of 3/5 .... You flip it and then switch the sign.

OpenStudy (icalibear):

Oh okay. So the equation for this line would be y = 3/5x + 4 1/5?

OpenStudy (anonymous):

Did you put (-2,3) into the equation of a line and solve?

OpenStudy (icalibear):

I just looked at the only multiple choice with 3/5 honestly, what is that step?

OpenStudy (anonymous):

(y - y's value) = m(x - x's value) m = slope and then solve Y-3=m(x-(-2)

OpenStudy (icalibear):

y - 3 = 3/5 (x - (-2)) ?

OpenStudy (anonymous):

Yep

OpenStudy (icalibear):

How do you simplify it?

OpenStudy (icalibear):

The 3 doesn't go there

OpenStudy (anonymous):

Y-3=3/5x+6/5 Y= 3/5x + 6/5 +3

OpenStudy (anonymous):

Y=3/5x+15/5

OpenStudy (icalibear):

ok. thanks!

OpenStudy (icalibear):

i got it :P

OpenStudy (anonymous):

Oops15/5 + 6/5

OpenStudy (anonymous):

Good job!

OpenStudy (icalibear):

I got it haha thank you :p

OpenStudy (icalibear):

thanks!